# Low-Rank Expectile Representations of a Data Matrix, with Application to Diurnal Heart Rates

### v2

@@ 3.3 Comparison of Optimization Algorithms
In our pursuit to minimize the loss function, we have selected three candidate gradient-based algorithms: Broyden–Fletcher–Goldfarb–Shanno algorithm (BFGS) [Gol70], Limited-Memory Broyden–Fletcher– Goldfarb–Shanno algorithm (LBFGS) [LN89], and Conjugate Gradient method (CG) [HS+52]. To evaluate the performance and efficiency of these algorithms, we utilized the previously generated simulated data to fit our model, conducting 10 trials for each algorithm. During these trials, we meticulously documented the loss function value at every iteration, alongside the final loss value obtained. Subsequently, we graphed both the raw and logarithmic loss function values, corresponding to a specific expectile $\tau$ and rank $k$.
The aggregated results, depicted in the figure 1 below, reveal distinctive patterns across hundreds of simulations. Notably, the BFGS algorithm exhibits a generally monotonic descent in loss value. Nonetheless, it occasionally encounters periods of stagnation, typically spanning 10-20 iterations, before resuming its downward trajectory. This intermittent stalling contributes to its considerably protracted computation time, typically exceeding that of LBFGS and CG by over a hundredfold. In contrast, the LBFGS and CG methods demonstrate higher convergence rates, with LBFGS outpacing CG. Despite their rapid convergence, both algorithms experience sporadic surges of instability in the loss function values at certain iterations. Ultimately, the final loss values attained by all three algorithms display a high degree of consistency.
@@
Figure 1: Snapshot of loss values for 3 algorithms when $\tau=0.3$ and $k=2$
@@ 3.5 Effect of misspecification of model rank
For instance, consistent with the previous discussion, the true rank of the simulated data’s $U$ and $V$ matrices is 2. We conducted simulations across different ranks $k=1,2,3,4$ with a constant $\tau=0.1$. The empirical outcomes, depicted in the subsequent figures, demonstrate that the final loss values for $k=1$ are approximately an order of magnitude greater than those obtained for $k=2,3,4$. Furthermore, while the loss values for $k=2,3,4$ exhibit remarkable consistency, the computational time required by the algorithm significantly increases with slight overfitting for $k=3$ and $4$, compared to the true rank of 2. We present the results for rank from 1 to 4 in the table 1 below.
@@ 4.2.3 Intraclass Correlation Coefficient Analysis
In this study, we calculated ICC values for each individual’s predictions of variables $C$ and $V$ separately to evaluate the consistency of our model’s predictions under different $\tau$ values. We constructed a new dataset by combining the prediction matrices of $C$ and $V$ with the corresponding participant IDs and detection dates. This setup allowed us to treat $Y$ as either $C$ or $V$ and $G$ as each individual participant, facilitating the computation of ICC values for different $\tau$ levels. The graph in 8 displays the ICC values with respect to different expectile values.

### v3

@@ 3.3 Comparison of Optimization Algorithms
In our pursuit to minimize the loss function, we have selected three candidate gradient-based algorithms: Broyden–Fletcher–Goldfarb–Shanno algorithm (BFGS) [Gol70], Limited-Memory Broyden–Fletcher– Goldfarb–Shanno algorithm (LBFGS) [LN89], and Conjugate Gradient method (CG) [HS+52]. To evaluate the performance and efficiency of these algorithms, we repetitively generated different simulated datasets as described above 100 times, and each time we conducted 50 initialization with specific $\tau=0.2$ and rank $k=3$. We recorded the average loss and time of each algorithm and find the algorithm with the smallest average loss and time respectively. Subsequently, we .
The aggregated results, depicted in the table 1 below, for each of the 100 different simulated datasets, the three algorithms have the minimum final loss almost evenly, and the differences in loss values among the three algorithms are extremely small, less than 0.00000001. While LBFGS outperforms significantly in the time of execution. It’s about 100 times faster than BFGS and 2-3 times faster than CG.
Algorithm Number of times with minimum final loss Number of times with minimum time of execution
BFGS 36 0
LBFGS 31 100
CG 33 0
Table 1: Algorithm comparison with 100 simulated datasets and 50 initialization of each dataset
@@ 3.5 Effect of misspecification of model rank
For instance, consistent with the previous discussion, the true rank of the simulated data’s $U$ and $V$ matrices is 2. We conducted simulations across different ranks $k=1,2,3,4$ with a constant $\tau=0.1$ for average of hundreds of trials. The empirical outcomes, depicted in the subsequent table, demonstrate that the final loss values for $k=1$ are approximately an order of magnitude greater than those obtained for $k=2,3,4$. Furthermore, while the loss values for $k=2,3,4$ exhibit remarkable consistency, the computational time required by the algorithm significantly increases with slight overfitting for $k=3$ and $4$, compared to the true rank of 2. We present the results for rank from 1 to 4 in the table 2 below.
@@ 4.2.3 Intraclass Correlation Coefficient Analysis
In this study, we calculated ICC values for each individual’s predictions of variables $C$ and $V$ separately to evaluate the consistency of our model’s predictions under different $\tau$ values. We constructed a new dataset by combining the prediction matrices of $C$ and $V$ with the corresponding participant IDs and detection dates. This setup allowed us to treat $Y$ as either $C$ or $V$ and $G$ as each individual participant, facilitating the computation of ICC values for different $\tau$ levels. The graph in 7 displays the ICC values with respect to different expectile values.